### v1

# thanks: Corresponding author

Abstract
In the present study, we propose a unique scheme to generate and control multiple flat bands in a decorated diamond chain by using a strain-induced proximity effect between the diagonal sites of each diamond plaquette. This is in complete contrast to the conventional diamond chain, in which the interplay between the lattice topology and an external magnetic flux leads to an extreme localization of the single-particle states, producing the flat bands in the energy spectrum. Such a strain-induced proximity effect will enable us to systematically control one of the diagonal hoppings in the decorated diamond chain, which will lead to the formation of both gapless and gapped flat bands in the energy spectrum. These gapless or gapped flat bands have been corroborated by the computation of the compact localized states amplitude distribution as well as the density of states of the system using a real space calculation. We have also shown that these flat bands are robust against the introduction of small amounts of random onsite disorder in the system. In addition to this, we have also classified the nontrivial topological properties of the system by calculating the winding numbers and edge states for the gapped energy spectrum. These findings could be easily realized experimentally using the laser-induced photonic lattice platforms.
I Introduction
Understanding the nature of exotic electronic states plays a pivotal role in building modern quantum technology-based devices. In quantum systems, the characteristics of the electronic states govern the quantum transport properties, which is a crucial component of quantum information processing and quantum communication. In general, there exist two different kinds of electronic states in quantum systems, viz., extended states and localized states. The first one is generally attributed to any wave-like phenomena in a periodic medium and known as the Bloch states Bloch-1929 , whereas the latter one describes the absence of diffusion of waves in a completely disordered or chaotic medium and manifests in the phenomena of Anderson localization Anderson-prb-1958 ; Ramakrishnan-prl-1979 . Later on, it was found that we can have exceptions to the above scenarios, where one can also have highly localized states even in a perfectly periodic system in the absence of any disorder. Such disorder-free localized states appearing in periodic lattice models are known as the flat-band (FB) states FB-review-Flach-2018 ; FB-review-Vicencio-2021 .
There are various reasons due to which these FB states emerge in different lattice models, such as the local topology of the lattice model Sutherland-prb-1986 , orbital frustration Trivedi-prb-2022 , influence of external magnetic flux Dias-prb-2011 ; Pal-prb-2018 ; Pal-jpcm-2025 , influence of electric field Flach-prb-2018 ; Flach-prr-2021 , hopping frustration Nat-Phys-2024 , and so on. At the FB energies, the kinetic energy of the particle is quenched, making it completely immobile, and hence it becomes highly localized. Due to this, the corresponding electronic wavefunction is caged over a couple of unit cells of the lattice geometry, and outside this region, it abruptly decays to zero. One can exactly construct the distribution of the wavefunction amplitudes for such localized states at the FB energies, which are known as the compact localized states (CLS) Ajith-prb-2017 ; Singular-FB-Rhim-2021 ; Pal-prb-2018 ; Pal-jpcm-2025 .
Figure 1: Schematic diagram of a quasi-one-dimensional decorated diamond lattice. The segment inside the dotted box indicates the unit cell of the lattice structure, consisting of four atomic sites. $t$ denotes the coupling along the periphery of each diamond (rhombic) structure, while $d_{H}$ and $d_{V}$ depict the diagonal couplings along the horizontal and vertical directions, respectively, inside a diamond structure. $\lambda$ represents the coupling in between two consecutive diamond structures.
Such FB states in the band structure of lattice models have attracted a lot of interest in recent times because they provide an excellent platform to explore various intriguing strongly correlated phenomena like the quantum Hall effect in the absence of a magnetic field Wen-prl-2011 ; Das-Sarma-prl-2011 ; Neupert-prl-2011 , FB superconductivity FB-superconductivity-1 ; FB-superconductivity-2 ; FB-superconductivity-3 ; FB-superconductivity-4 , inverse Anderson localization transition Inverse-Anderson-transition-1 ; Inverse-Anderson-transition-2 ; Inverse-Anderson-transition-3 , to name a few among others. These theoretical investigations on FB physics have been further accelerated by their experimental realization in the field of photonics Rodrigo-prl-2015 ; Sebabrata-prl-2015 ; Sebabrata-prl-2018 ; Rodrigo-prl-2022 . The rapid technological advancement in the fabrication and characterization techniques used in the field of photonics has empowered us to fabricate tailor-made photonic lattice structures with arbitrary lattice geometries and experimentally observe the localized FB states in those photonic lattice settings. Such techniques have enabled us to go beyond the regular lattice geometries and to explore FB states featuring in more complicated unusual lattice structures like super-honeycomb lattice Chen-Suepr-Honeycomb-AOM-2020 , super-Kagome lattice Chen-Super-Kagome-OL-2023 , decorated Lieb lattice Denz-Extended-Lieb-FB ; Bhatta-Pal-prb-2019 , and even in fractal-like lattices Pal-Saha-prb-2018 ; Hahafi-fractal-FB ; Chen-fractal-FB-1 ; Chen-fractal-FB-2 ; Biswas-fractal-FB . Of late, an exciting addition to the study of FB physics has been the observation of FB localization and topological edge states in topolectrical circuits Zhang-EC-FB-1 ; Zhang-EC-FB-2 ; Zhang-EC-FB-3 ; Flach-EC-FB .
The study of these FB states is not only important from the context of fundamental physics, but also it has various potential technological applications, such as quantum communication FB-quantum-commun , controlled optical transmission Sebabrata-prl-2018 ; Rodrigo-prl-2022 , image processing and precise optical measurement Sebabrata-prl-2015 , diffraction-free propagation of light Rodrigo-prl-2015 ; Sebabrata-prl-2015 , telecommunication and sensing devices Rodrigo-prl-2015 , and many more Flach-Nanophoton-2024 . Lattice models are widely used in physics for a variety of reasons, including approximating actual physical systems and offering useful settings for interesting theoretical investigation. Thus, simple lattice models hosting tunable FBs in their band structure are always significant for both technological applications and the investigation of novel fundamental physics. In the present study, we have considered one such simple quasi-one-dimensional decorated diamond lattice (also called a rhombic lattice) model (see Fig. 1) to explore the FB physics as well as the topological properties.
The conventional diamond lattice structures have been explored earlier in the context of extreme localization using an external magnetic flux. This phenomenon is known as the Aharonov-Bohm (AB) caging effect Vidal-prl-2000 ; Vidal-prl-1998 , where the particle is completely localized, giving rise to zero transport. In recent times, AB caging phenomena have been experimentally realized using photonic lattices Sebabrata-prl-2018 ; Rodrigo-prl-2022 and also using a synthetic lattice in the momentum space of ultracold atoms with tailored gauge fields Inverse-Anderson-transition-3 . Also, lately the phenomenon of AB caging has been theoretically explored for extended diamond chains Amrita-jpcm-2021 ; Sougata-PhysicaB-2024 and even for fractal structures built out of diamond-shaped structures Biplab-pssrrl-2025 . However, in all these studies, the extreme localization effect is induced by an external magnetic flux. In contrast to that, in the present study, we have examined a variant of the traditional diamond chain, where we have shown that FB localization can be tuned using a proximity effect. We have also demonstrated that the topological edge states in this lattice structure are caused by this proximity effect. In a very recent study, a similar antiferromagnetic diamond chain has been demonstrated to act as an efficient spin filter Debjani-AoP-2025 .
Our results and analysis follow next. The article is structured as follows: In Sec. II, we describe the construction and parameters of our lattice model and the tight-binding framework used to extract the results. The characterization of the proximity-induced FB states is demonstrated in Sec. III, which includes the construction of the CLS, computation of the density of states, and the effect of disorder on the FB states. In Sec. IV, we present the results related to the topological properties of the system, which includes calculation of the edge states and the winding number. Finally, we summarize the main findings of this study and its potential future scope in Sec. V.
II Description of the model
In this work, we have considered a quasi-one-dimensional decorated diamond lattice that grows along one direction only (see Fig. 1). The building block of this lattice model is a diamond (rhombic)-shaped structure that consists of four atomic sites, namely, $A$, $B$, $C$, and $D$, as shown in Fig. 1. The electronic properties of the single-electron states in this lattice model can be described by the tight-binding Hamiltonian written in the Wannier basis of the following form:
$\bm{H}=\sum_{n}\Big{[}\underbrace{\sum_{i}\varepsilon_{i}c_{n,i}^{\dagger}c_{n,i}}_{\textit{Onsite energy}}+\underbrace{\sum_{i,j}\Big{(}\mathcal{T}_{i,j}c_{n,i}^{\dagger}c_{n,j}+\textrm{h.c.}\Big{)}}_{\textit{Intra-cell hoppings}}+\underbrace{\sum_{\langle i,j\rangle}\Big{(}\lambda_{i,j}c_{n,i}^{\dagger}c_{n+1,j}+\textrm{h.c.}\Big{)}}_{\textit{Inter-cell hoppings}}\Big{]},$ (1)
where $\varepsilon_{i}$ represents the onsite energy of an electron to be placed at the $i$-th site ($i\in A,B,C,D$) in the $n$-th unit cell of the lattice. $\mathcal{T}_{i,j}$ represents the coupling (hopping amplitude) between the $i$-th and $j$-th sites within the same unit cell. $\mathcal{T}_{i,j}$ can take three possible values, viz., $t$ (when $i$ and $j$ are the two neighboring sites along the periphery of each diamond structure, i.e., in between a red and a cyan site), $d_{H}$ (when $i$ and $j$ represent the two neighboring sites along the horizontal diagonal of each diamond structure, i.e., in between two red sites), and $d_{V}$ (when $i$ and $j$ represent the two neighboring sites along the vertical diagonal of each diamond structure, i.e., in between two cyan sites) [see Fig. 1 for a better clarification]. $\lambda_{i,j}\equiv\lambda$ indicates the coupling between two neighboring sites $i$ and $j$ from the two neighboring unit cells, namely, $n$-th and $(n+1)$-th cells. $c_{n,i}^{\dagger}(c_{n,i})$ represents the creation (annihilation) operator for an electron at the $i$-th site in the $n$-th cell. We remark that, for our study, we have considered the spinless non-interacting electron formalism.
To compute the electronic band structure of the system, we transform the Hamiltonian in Eq. (1) from the Wannier basis description to the Bloch basis description using a discrete Fourier transform, which leads to
[…]
and $\bm{\tilde{c}}^{\dagger}_{k}=\left(\begin{matrix}c^{\dagger}_{k,A}&c^{\dagger}_{k,B}&c^{\dagger}_{k,C}&c^{\dagger}_{k,D}\end{matrix}\right)$. For the perfectly ordered system, we set the onsite potentials at all the sites $\varepsilon_{i}=0$. However, later we also investigate the effect of the random onsite disorder on the electronic properties of this lattice model, and in that case, we choose $\varepsilon_{i}$ randomly within an interval[Δ2,$[-\frac{\Delta}{2},\frac{\Delta}{2}]$, where $\Delta$ controls the strength of the random disorder. Throughout our calculation, we set the hopping parameters $\lambda=t=1$. We play with the hopping parameters $d_{H}$ and $d_{V}$, which can be appropriately controlled by the so-called proximity effect, to extract interesting FB features offered by this lattice model. This is elaborated in detail in Sec. III.
III Characterization of the proximity-induced flat bands
In this section, we discuss how to engender and control multiple FBs in this lattice structure. FBs appear in the band structure of certain lattice models or in real materials mainly as a result of destructive interference of the electron hoppings. In this lattice model, we realize such a destructive interference effect by controlling certain hopping parameters in the system with a proximity effect. This gives us an effective approach to manipulate multiple FBs in this lattice structure, as discussed below.
Figure 2: Electronic band structure of the decorated diamond lattice model. (a) A gapless FB appears at $E=0$ for $d_{H}=0$, (b) one gapless and one gapped FB appear at $E=0$ and $E=-2$, respectively, for $d_{H}=1$, (c) the FB at $E=0$ is gapped out for $d_{H}=1.5$, (d) two gapped FBs emerge at $E=0$ and $E=-1$, respectively, for $d_{H}=2$. We have set $d_{V}=0$ and $\lambda=t=1$ in all the four cases.
III.1 Electronic band structure
The Hamiltonian of a quantum system contains nearly all the crucial information about the system, making it an extremely useful tool for analyzing the electronic properties of the system. In order to extract the information about the electronic band structure of this lattice model, we diagonalize the Hamiltonian in Eq. (3). We keep the values of the hopping parameters $\lambda$ and $t$ unaltered, and tune the $d_{H}$ and $d_{V}$ couplings in the system to examine what happens to the band structure. First, we tune the value of $d_{H}$ (measured in units of $t$) by keeping $d_{V}=0$. In a real system, such an effect can be achieved either by distorting the lattice by means of applying precisely controlled pressure or using two different types of chemical species for the top-bottom and the horizontal atomic sites. It is found that the number of FBs and whether they are gapped or gapless with other dispersive bands can be suitably controlled by tuning the value of $d_{H}$. The results are shown in Fig. 2. Note that, the energy ($E$) is measured in units of $t$. It is worth mentioning that, this model always provides a very robust FB at the energy $E=0$ irrespective of the value of $d_{H}$. However, by tuning $d_{H}$, one can generate additional FB in the energy spectrum, as well as it can be gapped out from the other dispersive bands in the energy spectrum. The opening of the band gap in the energy spectrum is also useful to analyze the topological properties of this lattice model, as discussed in Sec. IV.
We note that, in this section, we have discussed the effect of tuning $d_{H}$ by keeping $d_{V}=0$. This makes sense, because as we increase the value of $d_{H}$, the coupling between the horizontal diagonal sites increases, which effectively means that the coupling between the vertical diagonal sites will vanish. However, it is also interesting to look into the opposite scenario, where we make the value of $d_{V}$ nonzero by keeping $d_{H}=0$. This case is discussed in the Appendix A.
III.2 Compact localized states
Figure 3: The real-space wave function amplitude distribution corresponding to the flat bands at the energies (a) $E=0$, (b) $E=-1$, and (c) $E=-2$. Zero wave function amplitudes are denoted by grey-shaded circles, while the nonzero wave function amplitudes are marked by green (for positive amplitudes) and red (for negative amplitudes) circles, respectively. The system parameters are chosen to be the same as in Fig. 2.
The appearance of the FBs in the energy spectrum elaborated in subsection III.1 is a momentum ($k$)-space description. In the present subsection, we will discuss how one can classify such FB states using a real space description. It is in general true that, corresponding to the FB energies in the electronic band structure of a lattice model, it is always possible to construct a special kind of eigenstate in the real space in which we can have nonzero wave function amplitude distribution only over a few unit cells of the lattice, and they are connected with each other with strictly zero wave function amplitudes. Such real-space configurations of the FB states are known as the compact localized states. The construction of the CLS can be imagined to act as a quantum prison, in which the electrons are completely localized at the FB energies even in the absence of any disorder in the system.
For the decorated diamond lattice model, the construction of the CLS corresponding to the various FB energies is exhibited in Fig. 3. We observe that for the FB at $E=0$, the CLS spans over only one unit cell, whereas for the FBs at $E=-1$ and $E=-2$, the CLS are spread over two unit cells. Such spread of the CLS is often connected to a boost of the quantum metric, which is the real part of the quantum geometric tensor, and may be useful in the study of unconventional FB superconductivity Quantum-metric-prb-2022 . The highly localized character of the eigenstates corresponding to the CLS will be more clear to us from the analysis of the density of states of the system, which is discussed in the following subsection.
III.3 The density of states
Figure 4: Variation of the average density of states (ADOS) as a function of the energy ($E$) for the decorated diamond chain. The total number of unit cells considered here is $100$ (i.e., $\mathcal{N}=400$ sites). The plots are done for different values of the horizontal diagonal coupling $d_{H}$ (measured in units of $t$), viz., (a) $d_{H}=0$, (b) $d_{H}=1$, (c) $d_{H}=1.5$, and (d) $d_{H}=2$. The other parameters are chosen as $d_{V}=0$ and $\lambda=t=1$ in all four cases. The value of the small imaginary part $\eta$ is taken to be equal to 0.005.
The highly localized character of the FB states in a lattice model can be distinctly identified by looking at its density of states spectrum. One can easily compute the average density of states (ADOS) of a finite-size system by using the Green’s function formalism given by,
$\rho(E)=-\dfrac{1}{\mathcal{N}\pi}\textrm{Im}\Big{[}\textrm{Tr}\left[\bm{\mathcal{G}}(E)\right]\Big{]},$ (4)
where $\bm{\mathcal{G}}(E)=\big{[}(E+i\eta){\bm{I}}-\bm{H}\big{]}^{-1}$ (with $\eta\rightarrow 0^{+}$), ${\bm{I}}$ is an identity matrix of $\mathcal{N}\times\mathcal{N}$ dimension, $\mathcal{N}$ being the total number of sites in the lattice, and ‘Tr’ denotes the trace of the Green’s function matrix $\bm{\mathcal{G}}$. By using the above formula, we have computed the ADOS for a finite-size decorated diamond chain with $100$ unit cells (i.e., $\mathcal{N}=400$ sites). The results are displayed in Fig. 4. From Fig. 4, we can clearly identify the appearance of the sharp peaks in the ADOS spectrum corresponding to the highly localized states exactly at the FB energies. The other relatively smaller picks indicate the band edges. From the ADOS spectrum, one can clearly distinguish whether the FBs are gapped or gapless from the other dispersive continuum bands. We note that the ADOS spectrum exactly corroborates the information we have obtained about the FBs in Fig. 2.
III.4 The robustness of the FBs against disorder
Figure 5: Variation of the average density of states (ADOS) as a function of the energy ($E$) for the decorated diamond chain by incorporating a random onsite (diagonal) disorder of strength $\Delta=0.2$ in the system. All other parameters are kept unaltered as in Fig. 4.
It is very important to examine the robustness of these FB states against the small amounts of disorder in the system. In this subsection, we discuss this aspect. In general, when a flat band is separated from the other dispersive bands in the spectrum by a gap, the compact localized states are robust against the disorder, retaining their strong localization character under the finite disorder strength. However, when the band gap vanishes, the effective disorder potential is enhanced. In that scenario, the compact localized states are destroyed, and they acquire large localization lengths with different power law scalings as compared to the Anderson localization length.
For our lattice model, we introduce a random diagonal disorder in the onsite potential distribution of the system $\varepsilon_{i}\sim[-\frac{\Delta}{2},\frac{\Delta}{2}]$, where $\Delta$ denotes the disorder strength. We find that, for a relatively small disorder strength $\Delta=0.2$ (measured in units of $t$), the highly localized FB states in the ADOS spectrum remain intact as depicted in Fig. 5. However, as we increase the strength of the onsite disorder $\Delta$ to higher values, the highly localized FB states collapse, and we can have a localization-to-delocalization transition in the system known as the inverse Anderson transition Inverse-Anderson-transition-1 ; Inverse-Anderson-transition-2 ; Inverse-Anderson-transition-3 .
IV Topological properties of the model
Figure 6: (a) Energy eigenspectra of the lattice model as a function of the horizontal diagonal coupling $d_{H}$ to mark the emergence of the edge states by tuning $d_{H}$. The edge states are two-fold degenerate and appear after a critical value of $d_{H}$ at the energy $E=-\sqrt{2}$. (b) Distribution of the energy eigenspectra for $d_{H}=t$ with an open boundary condition for $50$ unit cells. The in-gap degenerate edge states are marked with an encircled red dot at the energy eigenvalue $E=-\sqrt{2}$. (c, d) Amplitude distributions of one pair of in-gap edge states with energy $E=-\sqrt{2}$ for $50$ unit cells with an open boundary condition. The distribution for one state is confined to the left edge (shown in violet color), while the other one is strictly localized at the right edge (shown in green color). The other parameters are fixed at $\lambda=t=1$ and $d_{V}=0$.
In the present section, we present the results related to the topological properties offered by the decorated diamond lattice model. Comprehending the topological characteristics of a lattice model often has important ramifications for emerging quantum technologies. For our lattice model, the topological nature of the bands is characterized by the Zak phase or, equivalently, by the winding number. This topological invariant corresponds to the Berry phase berry-1984 , which can be expressed in terms of the $n$-th Bloch eigenstates as
[…]
The Zak phase is related to the winding number by $\mathcal{Z}=\pi\nu$. Experimentally, the topological signature of a system has been demonstrated through the measurement of the winding number in analogous photonic networks exp-winding-2018 ; exp-winding-2020 . In our analysis, we compute the winding number using the gauge-invariant Wilson loop method zak-fukui-2005 . According to Eq. (5), the integration is converted into the summation over the entire Brillouin Zone by dividing it into $N$ number of similar discrete sections, which ensures that each interval in the wave vector will be $\delta k=2\pi/Na$ with lattice constant $a$. For our computation, we consider $N=401$, which provides a sufficiently accurate approximation of the integration. Within the Wilson loop framework, for a non-degenerate $n$-th band, the winding number is given by zak-fukui-2005 ; zak-marzari-1997 ,
$\mathcal{\nu}=-\dfrac{1}{\pi}\textrm{Im}\bigg{[}\log\Big{(}\prod_{k_{j}}\langle u_{k_{j},n}|u_{k_{j+1},n}\rangle\Big{)}\bigg{]}.$ (6)
Till now, we have discussed the quasi-one-dimensional decorated diamond lattice with periodic boundary conditions. Now, we focus on the scenario with open boundary condition to explain the emergence of topological nontrivial features for this lattice model as exhibited in Fig. 6. Fig. 6(a) explains the energy spectrum of the system as a function of the horizontal diagonal coupling parameter $d_{H}$. Beyond a critical value of $d_{H}$, we observe the emergence of in-gap edge states, which persist over a finite range of $d_{H}$. Within this regime, the bulk bands become gapped after a band-closing transition, indicating a topological phase transition into a nontrivial phase. These edge states disappear upon imposing periodic boundary conditions, confirming their boundary-localized nature. In particular, Fig. 6(a) reveals that the edge states are pinned at energy $E=-\sqrt{2}$ for a specific range of $d_{H}$. Throughout these calculations, we fix other parameters $\lambda=t=1$ and $d_{V}=0$.
Figure 7: The variation of the winding number as a function of $d_{H}$ for the lowest band, which clearly highlights the topologically nontrivial regime with winding number $-1$.
To further explore this, we consider $d_{H}=t$ and plot the eigenvalues with respect to the site index for a system size with $50$ unit cells in Fig. 6(b). The degenerate edge states are clearly identified and marked with an encircled red dot, showing a two-fold degeneracy. As depicted in Fig. 6(c) and (d), one of these degenerate edge states is localized at the left edge of the system, while its partner is confined to the right edge of the system. These states are exponentially localized near the edges and decay very rapidly into the bulk, which is a hallmark of the topologically protected modes. To characterize the topological nature of this phase, we compute the winding number as a function of $d_{H}$, which yields integer values within the nontrivial topological zone. This quantized invariant confirms the presence of topologically protected edge states through the principle of bulk-boundary correspondence. The winding number distribution, shown schematically, clearly distinguishes the topologically trivial and nontrivial phases, with nonzero integer values of winding numbers indicating topological behavior (see Fig. 7). We have analyzed the variation of the winding number specifically for the lowest occupied gapped band to characterize the evolution of topological features in the system. The winding number is well-defined only when the corresponding band structure remains gapped. Notably, edge states emerge beyond a critical value of $d_{H}$, where a gap opens up between the lowest and the second bands. In this nontrivial topological regime, the winding number for the lowest band takes the value $-1$, while that for the second band is $1$. In contrast, the winding numbers for the third and the highest bands are zero, indicating their topologically trivial nature. It is worthwhile to mention that, the winding number confirms the number of edge states. These edge states are non-chiral in nature, as they are not only restricted to a single edge of the lattice; instead, each member of the degenerate pair is localized at the opposite edges of the system.
V Concluding remarks
In conclusion, we have studied the emergence of multiple numbers of nondispersive, completely flat bands in a quasi-one-dimensional decorated diamond chain model. It has been shown that, these flat bands can be controlled by a proximity effect, which regulates the horizontal diagonal coupling strength in each diamond unit cell. The appearance of these flat band states in the system is substantiated by the construction of the compact localized states and the computation of the average density of states. It has also been verified that, these flat bands are robust against the introduction of small amounts of random disorder in the onsite potential distribution of the system. Finally, we have studied the topological features of this lattice model by calculating the topologically protected edge states and the winding numbers using the Wilson loop approach. These interesting findings for such a simple quantum geometry make it a prospective candidate to be used to create a novel, cutting-edge device based on the quantum technologies in the future. Additionally, because of the simplicity of the proposed lattice model, it can be easily built and accessed in a real-life experiment using the laser-induced photonic lattice settings.
Acknowledgements.
V.T. acknowledges the Govt. of India for providing funding through a NFST research fellowship. B.P. would like to thank Prof. Georges Bouzerar and Maxime Thumin from the Néel Institute for useful discussions on the effect of disorder on the flat bands.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Effect of vertical diagonal coupling $d_{V}\neq 0$
Figure A.1: Electronic band structure of the decorated diamond lattice model with nonzero values of $d_{V}$. For nonzero finite values of $d_{V}$, the FB appears at $E=-d_{V}$. (a) For $d_{V}=1$, a gapless FB appears at $E=-1$ and (b) For $d_{V}=2.5$, a gapped FB appears at $E=-2.5$. The other parameters are set to be $d_{H}=0$ and $\lambda=t=1$.
In the main text of the manuscript, we have focused on the effect of the horizontal diagonal coupling $d_{H}$ on the band structure of this lattice model by keeping the vertical diagonal coupling $d_{V}=0$. However, it is a pertinent question to ask what happens for $d_{V}\neq 0$. This question is addressed in this appendix. We find that, for $d_{V}\neq 0$ (keeping $d_{H}=0$), a single FB always appears in the band structure at the energy $E=-d_{V}$ (measured in units of $t$). Two sample cases are shown in Fig. A.1. This result is also very interesting, as it gives us the flexibility to set the FB precisely at any desired finite value of the energy by tuning the vertical diagonal coupling $d_{V}$ in the system.